Solve the system of equations by using the substitution method.
{2x – 3y=14
{y=x - 5

Found 2 solutions by mananth, jorel1380:
Answer by mananth(16946)

2 x -3 y = 14 .............1
1 y = -5 + 1 x ---------2
/ 1 =
y = -5 + 1 x
Substitute y in (1)
2 x -3 ( -5 + 1 x ) = 14
2 x + 15 + -3 x = 14
2 x -3 x = 14 -15
-1 x = -1
/ -1
x= 1
Plug the value of x in (1)
2 x -3 y = 14
2 * 1 -3 y = 14
2 -3 y = 14
-3 y = 12
/ -3
y= -4
m.ananth@hotmail.ca

Answer by jorel1380(3719)

2x – 3y=14
y=x - 5
2x-3(x-5)=14
2x-3x+15=14
-x=-1
x=1
y=-4..